A flower bed is in the shape of a triangle. One side is twice the length of the shortest side,
and the third side is 10 feet more than the length of the shortest side. The perimeter is 38 feet.
Find the lengths of all three sides.

Answer by ewatrrr(24785)

Hi,
Let x represent the shortest side
.
Question states
x + 2x + (x + 10) = 38 ft
.
simplify and solve
4x + 10 = 38ft
4x = 28
x = 7 ft
2x = 14 ft
(x+10) = 17ft
.
check your answer
7ft +14ft + 17ft = 38ft